OpenStudy (anonymous):

for the equation |-1 + x| = 5 would the answer be four or six? I can't figure it out.

OpenStudy (welshfella):

the vertical lines around the left side of the equation mean the absolute value - that means the positive value is taken even if the inside part is negative

OpenStudy (welshfella):

so we write 2 equations -1 + x = 5 and -1 + x = -5 solve these 2 for x

OpenStudy (anonymous):

@welshfella I know! I got 4 from that but my friend whose on the same lesson said it was 6 so i'm slightly confused

OpenStudy (welshfella):

there are 2 solutions -1 + x = 5 gives x = 6 and -x + 1 = -5 gives x = -4

OpenStudy (welshfella):

sorry the last ones is wron -x + 1 = -5 -x = -4 x = 4

OpenStudy (welshfella):

so the 2 solutions are 4 and 6

OpenStudy (rizags):

i believe its -4, 6

OpenStudy (anonymous):

should i put both as an answer then?

myininaya (myininaya):

yes -1+x=-5 should give x=-5+1

OpenStudy (rizags):

both are correct

OpenStudy (welshfella):

yes - i made a typo when i wrote the second equation down

myininaya (myininaya):

|-1+x|=5 gives -1+x=5 or -1+x=-5

OpenStudy (welshfella):

yes i wrote -x + 1 = -5 instead of -1 + x = -5

OpenStudy (welshfella):

-4 , 6 is correct